Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by swerve » Fri Nov 13, 2020 10:20 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?

A. 1/8
B. 1/4
C. 1/2
D. 3/4
E. 7/8

The OA is D

Source: Princeton Review
Join the discussion
Source: — Problem Solving |

Re: Probability

by Scott@TargetTestPrep » Mon Nov 16, 2020 12:25 pm
swerve wrote: ↑
Fri Nov 13, 2020 10:20 am
If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?

A. 1/8
B. 1/4
C. 1/2
D. 3/4
E. 7/8

The OA is D

Solution:

The phrases “at least one even number” and “at least one odd number” translate as “two even numbers and one odd number” or “one even number and two odd numbers,”. The probability of getting the former is 1/2 x 1/2 x 1/2 x 3C2 = 1/8 x 3 = 3/8 (notice that 3C2 = 3 is the number of ways of arranging two even numbers and one odd number). Similarly, the probability of getting the latter is also 3/8. Therefore, the overall probability is 3/8 + 3/8 = 6/8 = 3/4.

Alternate Solution:

When a six sided die is rolled three times, the total number of possible outcomes is 6 x 6 x 6 = 216.

Notice that if “getting at least one even number and at least one odd number” is not satisfied, then the outcome consists of either all even or all odd numbers. Of the 216 possible outcomes, 3 x 3 x 3 = 27 of them consist of all even numbers and by the same reasoning, another 27 of them consist of all odd numbers. Thus, 27 + 27 = 54 of the outcomes do not meet the requirement of containing at least one even number and at least one odd number. It follows that 216 - 54 = 162 of the outcomes meet this requirement and hence, the probability is 162/216 = 3/4.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion